OpenStudy (anonymous):

A function is given below. Determine the average rate of change of the function between x = 3 and x = 3 + h. g(x) = -7/x How do i solve this? I don't seem to be getting the right answer...

OpenStudy (campbell_st):

find g(3) and g( 3 +h) average rate of change is \[\frac{g(3 + h) - g(3)}{ 3+h - 2}\] substitute and evaluate

OpenStudy (anonymous):

What happened to the -7 though?

OpenStudy (campbell_st):

oops denominator should be 3 + h - 3

OpenStudy (campbell_st):

\[g(3) = -\frac{7}{3}\] \[g(3 +h) = -\frac{7}{ 3 + h}\]

OpenStudy (campbell_st):

so you have average rate of change \[\frac{-\frac{7}{3 +h} - -\frac{7}{3}}{ 3 +h - 3}\]

OpenStudy (anonymous):

Should I get the -7/3+h - -7/3 under the same denominator first or o I start canceling as is?

OpenStudy (campbell_st):

thats correct... now simplify.... you need a common denominator in the numerator

OpenStudy (anonymous):

7h/9+3h for the denominator in the numerator?

OpenStudy (anonymous):

or is this wrong?

OpenStudy (campbell_st):

thats correct, now just divide by h

OpenStudy (anonymous):

I get 7/9h+3 as my final answer because I multiplied by the inverse?

OpenStudy (campbell_st):

thats what I got... well done

OpenStudy (anonymous):

The computer said its wrong

OpenStudy (anonymous):

I only get one more attempt before it locks me out of a quiz and gives me a 0 on it...

OpenStudy (campbell_st):

well its my solution.... may be just in the way the answer is written it could be \[\frac{7}{3(3 + h)}\] if you're in doubt... use wolframalpha...

OpenStudy (anonymous):

Ok, thanks for your help I appreciate it!